If the market price is 30,inverse market demand is specified by P=120-0.18Q,the industry supply is consistent with P=10+0.04Q,what is the consumer surplus?

1
Expert's answer
2022-04-07T13:08:09-0400

The equilibrium quantity can be found from the equation 120-0.18Q=10+0.04Q. Q=500. From the equation of demand we can find the price level when Q=0. P=120-0.18×0=120. The consumer surplus is 1/2×(120-30)×500=22500